A disk-shaped merry-go-round of radius 2.43 m and mass 185 kg rotates freely with an angular speed of 0.641 rev/s . A 62.4 kg person running tangential to the rim of the merry-go-round at 3.11 m/s jumps onto its rim and holds on. Before jumping on the merry-go-round, the person was moving in the same direction as the merry-go-round's rim. Part A What is the final angular speed of the merry-go-round?
Answer #1

0.641 rev/s = 4.03 rad/s

I = 1/2 m r2 = 1/2 * 185 * 2.432 = 546.2 kg.m2

Li = I w + M r v = (546.2 * 4.03) + (62.4 * 2.43 * 3.11) = 2673 kg.m2/s

Lf = I + M r2 = (546.2 ) + (62.4 * 2.432) = 915 kg.m2/s

final angular speed = 2673 /  915

final angular speed = 2.92 rad/s